# Quantum Many-Body Scars for Arbitrary Integer Spin in $2+1$D Abelian Gauge Theories

### v1

@@ Abstract
The existence of Quantum Many-Body Scars, high-energy eigenstates that evade the Eigenstate Thermalization Hypothesis, has been established across different quantum many-body systems, including gauge theories corresponding to spin-1/2 Quantum Link Models. We systematically identify scars for pure gauge theories with arbitrarily large integer spin $S$ in $2+1$D, concretely for Truncated Link Models, where the electric field is restricted to $12S+$ 1 states per link. Through an explicit analytic construction, we show that the presence of scars is widespread in $2+1$D gauge theories for arbitrary integer spin. We confirm these findings numerically for small truncated spin and $S=1$ Quantum Link Models. The proposed analytic construction establishes the presence of scars far beyond volumes and spins that can be probed with existing numerical methods.
@@
The effect of QMBS was first observed in a Rydberg-atom quantum simulator, marked by persistent revivals for particular initial states, in contrast to the vast majority of other high-energy initial states [10, 11]. In parallel with the exact construction of highly excited eigenstates in a non-integrable model [12, 13], the experiment has sparked intense research across different systems, where a weak breaking of the ETH takes place [14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. In this case, the presence of few anomalous eigenstates, can leave an imprint on thermalization. Among their special features is that they are characterized by atypically low entanglement entropy, compared with other states arbitrarily close in the spectrum. This is in stark contrast to what is expected from the ETH. In particular, QMBS may exhibit an area law. This means that the entanglement between a subsystem and its complement is proportional to the area of the boundary that divides them. In turn, highly excited states are expected to have an entanglement that grows linearly with the volume of the subsystem.
@@
The interplay between gauge symmetry and scarring phenomena is not fully understood. Gauge theories naturally lead to constraints, and specific constraints can give rise to QMBS, as observed in the PXP model [10, 24, 25]. Gauge theories also feature local symmetries, which segment the Hamiltonian into distinct sectors that do not mix under time evolution. Supplied by a condition on physical states, this leads to local constraints in the form of Gauss’ law. Beyond the PXP model, QMBS have also been observed in other gauge theories, such as the Abelian case in the presence of matter in $1+1$D [26], pure gauge in $2+1$D [27, 28, 29] and non-Abelian in $2+1$D [30]. So far, the research on QMBS in Abelian pure gauge theories has exclusively focused on spin-$1/2$ Quantum Link Models (QLM) [27, 28, 29], which also exhibit other interesting phenomena such as crystalline confining phases [31]. This Letter demonstrates, for the first time, the existence of QMBS in link models with arbitrary integer spins.
@@
Addressing questions related to real-time dynamics falls outside the capabilities of conventional lattice gauge theory methods, as Monte Carlo simulations face inefficiencies due to severe sign problems associated with real-time processes (see e.g. [32, 33]). Sign problems have motivated the development of the field of quantum simulations of gauge theories [34, 35, 36, 37, 38, 39]. Formulations like QLMs are essential for preserving gauge symmetry while ensuring a finite Hilbert space for a finite lattice volume. Recovering the theory in the continuum limit may take different routes [40, 41, 42, 43, 44, 45]. No matter the path, the road ahead is long until we can make decisive contributions to theories like Quantum Chromodynamics (QCD). Hence, it is imperative to pinpoint intriguing phenomena that can drive experiments along this path. The study of QMBS is a meritorious example as it is present in simple gauge theories and probes fundamental aspects of quantum many-body theory.
@@
We show that it is also possible to construct such states for arbitrary spin truncated link models. The number of these low entropy states grows exponentially with volume. By choosing appropriately the potential term of the Hamiltonian, different linear combinations can be isolated as low entropy eigenstates. To establish these states as scars, one should prove that the model in question is non-integrable. Since we can construct these states for a wide region of parameter space, we expect this to be the case.
@@
Figure 1: Left: Example of a tiling $T$ of the 2D lattice, consisting of partitioning it into $2\times 1$ and $1\times 2$ tiles, to a total of 8 tiles. Right: Representation of a scar state based on an alternative tiling and in the dual representation. An independent sum over every $k_{i}$ with a weight of $\prod_{i}\left(-1\right)^{k_{i}}$ is assumed. See Eq. (19).
@@
In the dual representation, the state (19) in a $2\times 1$ plaquette system is
@@
All states that contribute in (19) have a height variable representation where there is one state out of the sum (7) on all tiles of the tiling $T$. The state sums over all $(S+1)^{\left|T\right|}$ combinations of these states. A pictorial representation of this tiling is shown on the right-hand side of Fig. 1. This tiling structure is similar to the lego and sub-lattice scars constructed for the $S=1/2$ case [28, 29], and allows for the construction of an exponential number of scars for arbitrary truncated link models. A formal definition of this tiling, in the form of a tiling product, can be found in the supplementary material [46]. There, we also describe a more general framework to construct more low-entropy zero-modes in TLMs.
@@
The scar has the form (19), just as the other scars we derived. However, the two plaquettes that make up tiles do not share a link, but neighbor each other diagonally in a pattern depicted in Fig. 4. This is still a zero-mode since each plaquette has one neighbor with height ±$\pm 1$ in all configurations and may therefore never take height $\mp 1$ itself. The state is an eigenstate of (10) because the number of neighbors with height ±$\pm 1$ is always the same for both plaquettes of each tile. This means that in all configurations, half of the links have values ±$\pm 1$ and the others are 00. Establishing the presence of these scars beyond $S=1$ is beyond the scope of this work. Systems with larger spins are less accessible through ED beyond single-leg ladders and we did not observe any of these scars for $S\leq 2$ in this case.
@@
Note.—During the final stages of our manuscript, we became aware of another work [52] on quantum many-body scarring in a 2+1D U(1) gauge theory with dynamical matter. This work will appear in the same arXiv listing as ours.

### v2

@@ Abstract
The existence of Quantum Many-Body Scars has been established across different quantum many-body systems, including spin-1/2 Quantum Link Models. We systematically identify scars for pure gauge theories with arbitrarily large integer spin $S$ in $2+1$D, concretely for Truncated Link Models, where the electric field is truncated rather than replaced by spins. Through an explicit analytic construction, we show that the presence of scars is widespread. We confirm these findings numerically for small truncated spin and $S=1$ Quantum Link Models. The proposed analytic construction establishes the presence of scars far beyond volumes and spins that can be probed with existing numerical methods.
@@
The effect of QMBS was first observed in a Rydberg-atom quantum simulator, marked by persistent revivals for particular initial states, in contrast to the vast majority of other high-energy initial states [10, 11]. In parallel with the exact construction of highly excited eigenstates in a non-integrable model [12, 13], the experiment has sparked intense research across different systems, where a weak breaking of the ETH takes place [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. In this case, the presence of few anomalous eigenstates, can leave an imprint on thermalization. Among their special features is that they are characterized by atypically low entanglement entropy, compared with other states arbitrarily close in the spectrum. This is in stark contrast to what is expected from the ETH. In particular, QMBS may exhibit an area law. This means that the entanglement between a subsystem and its complement is proportional to the area of the boundary that divides them. In turn, highly excited states are expected to have an entanglement that grows linearly with the volume of the subsystem.
@@
The interplay between gauge symmetry and scarring phenomena is not fully understood. Gauge theories naturally lead to constraints, and specific constraints can give rise to QMBS, as observed in the PXP model [10, 27, 28]. Gauge theories also feature local symmetries, which segment the Hamiltonian into distinct sectors that do not mix under time evolution. Supplied by a condition on physical states, this leads to local constraints in the form of Gauss’ law. Beyond the PXP model, QMBS have also been observed in other gauge theories, such as the Abelian case in the presence of matter in $1+1$D [29], pure gauge in $2+1$D [30, 31, 32] and non-Abelian in $2+1$D [33]. So far, the research on QMBS in Abelian pure gauge theories has exclusively focused on spin-$1/2$ Quantum Link Models (QLM) [30, 31, 32], which also exhibit other interesting phenomena such as crystalline confining phases or confining strings with fractionalized electric flux stands [34, 35, 36, 37]. This Letter demonstrates, for the first time, the existence of QMBS in link models with arbitrary integer spins.
@@
Addressing questions related to real-time dynamics falls outside the capabilities of conventional lattice gauge theory methods, as Monte Carlo simulations face inefficiencies due to severe sign problems associated with real-time processes (see e.g. [38, 39]). Sign problems have motivated the development of the field of quantum simulations of gauge theories [40, 41, 42, 43, 44, 45]. Formulations like QLMs are essential for preserving gauge symmetry while ensuring a finite Hilbert space for a finite lattice volume. Recovering the theory in the continuum limit may take different routes [46, 47, 48, 49, 50, 51, 52]. No matter the path, the road ahead is long until we can make decisive contributions to theories like Quantum Chromodynamics (QCD). Hence, it is imperative to pinpoint intriguing phenomena that can drive experiments along this path. The study of QMBS is a meritorious example as it is present in simple gauge theories and probes fundamental aspects of quantum many-body theory.
@@
We show that it is also possible to construct such states for arbitrary spin truncated link models. The number of these low entropy states grows exponentially with volume. By choosing appropriately the potential term of the Hamiltonian, different linear combinations can be isolated as low entropy eigenstates. To establish these states as scars, one should show that the model in question is non-integrable. It is expected that this is the case for pure gauge theories in (2+1)-d, and we further demonstrate this for a single-leg ladder with $S=1$ in the Supplementary Material [53].
@@
Figure 1: Left: Example of a tiling $T$ of the 2D lattice, consisting of partitioning it into $2\times 1$ and $1\times 2$ tiles, to a total of 8 tiles. Right: Representation of a scar state based on an alternative tiling and in the dual representation. An independent sum over every $k_{i}$ with a weight of $\prod_{i}\left(-1\right)^{k_{i}}$ is assumed. See Eq. (Quantum Many-Body Scars for Arbitrary Integer Spin in $2+1$D Abelian Gauge Theories).
@@
In the dual representation, the state (Quantum Many-Body Scars for Arbitrary Integer Spin in $2+1$D Abelian Gauge Theories) in a $2\times 1$ plaquette system is
@@
All states that contribute in (Quantum Many-Body Scars for Arbitrary Integer Spin in $2+1$D Abelian Gauge Theories) have a height variable representation where there is one state out of the sum (7) on all tiles of the tiling $T$. The state sums over all $(S+1)^{\left|T\right|}$ combinations of these states. A pictorial representation of this tiling is shown on the right-hand side of Fig. 1. This tiling structure is similar to the lego and sub-lattice scars constructed for the $S=1/2$ case [31, 32], and allows for the construction of an exponential number of scars for arbitrary truncated link models. A formal definition of this tiling, in the form of a tiling product, can be found in the Supplementary Material [53]. There, we also describe a more general framework to construct more low-entropy zero-modes in TLMs.
@@
The scar has the form (Quantum Many-Body Scars for Arbitrary Integer Spin in $2+1$D Abelian Gauge Theories), just as the other scars we derived. However, the two plaquettes that make up tiles do not share a link, but neighbor each other diagonally in a pattern depicted in Fig. 4. This is still a zero-mode since each plaquette has one neighbor with height ±$\pm 1$ in all configurations and may therefore never take height $\mp 1$ itself. The state is an eigenstate of (10) because the number of neighbors with height ±$\pm 1$ is always the same for both plaquettes of each tile. This means that in all configurations, half of the links have values ±$\pm 1$ and the others are 00. Establishing the presence of these scars beyond $S=1$ is beyond the scope of this work. Systems with larger spins are less accessible through ED beyond single-leg ladders and we did not observe any of these scars for $S\leq 2$ in this case.
@@
Note.—During the final stages of our manuscript, we became aware of another work [59] on quantum many-body scarring in a 2+1D U(1) gauge theory with dynamical matter.